OpenStudy (anonymous):

Which logarithmic graph can be used to approximate the value of y in the equation 2y = 5?

OpenStudy (irishboy123):

\(y = 2.5\) or do you mean \(2^y = 5\) ??!!

OpenStudy (anonymous):

2^y=5. the second thing you said

OpenStudy (phi):

if you start with 2^y= x then y log(2)= log(x) y= log(x)/log(2) the graph of this should look something like